# High-flux cold lithium-6 and rubidium-87 atoms from compact two-dimensional magneto-optical traps

### v1

@@ III.1 Evaluating the Li short-distance Zeeman slower
where $R$ is the loading rate, $\Gamma=1/\tau$ is the loss rate coefficient arising from background collisions, and $\beta$ is the two-body collisional loss coefficient. Solving this differential equation with the initial condition, $N_{0}=0$ yields the expression of the MOT dynamics (assuming isotropic density distribution)[47]
@@
The Zeeman slower significantly boosts the atomic flux, thereby enhancing both the loading rate and the saturation atom number. Specifically, the loading rate increases by a factor of 44 to $R^{\mathrm{Li}}_{\mathrm{zs}}=6.639(25)\times 10^{9}~\mathrm{atoms/s}$ at a moderate oven temperature of $T_{\mathrm{Li}}=372~^{\circ}\mathrm{C}$. This temperature is $80~^{\circ}\mathrm{C}$ lower (23 times lower vapor pressure) than typical $^{6}\mathrm{Li}$ experiments using traditional Zeeman slowing techniques [17], which also contributes to prolonging the lifetime of the $^{6}\mathrm{Li}$ source.
@@ III.2 Li flux dependence
Figure 6: Optimization of the Zeeman slowing parameters. (a) 3D MOT loading rate as a function of Li oven temperature. The solid line represents the theoretical Li vapor pressure in the liquid phase [38]. (b) Loading rate dependence on the repumper-to-cooler intensity ratio. (c) Loading rate dependence on the Zeeman beam power. Measurements in panels (b) and (c) were conducted at a fixed oven temperature of $335\,^{\circ}\mathrm{C}$.
@@ IV Summary and Conclusion
During the preparation of this manuscript, we became aware of a compact high-flux ytterbium-rubidium Zeeman slower [48].
@@ Appendix A Differential pumping channels design
It is essential to carefully design the dimensions of the differential pumping channels (DP) between the chambers to maintain a sufficient pressure ratio while not blocking the atomic flux more than necessary. The atomic beam divergence is given as its transverse-to-longitudinal velocity ratio, $\zeta=v_{\mathrm{tr}}/v_{x}$. The first is a result of the finite transverse temperature after the 2D MOT stage, while the latter can be tuned via the push beam power. In general, it is advantageous to reduce the distances from 2D MOTs to the science cell as much as possible, resulting in a larger opening angle of the science cell and thus an increased capturable fraction of atoms. The DP dimensions are then chosen to not clip the atomic beam significantly more than the science glass cell, which is shown schematically in Figure 9. The DPs consist of two sections in series (with diameters $d_{1}$ and $d_{2}$ as well as lengths $l_{1}$ and $l_{2}$), as it is easier to manufacture compared to one conical tube (as used, e.g., in [30, 49]). Additionally, the Rb DP is 45 polished on the surface facing the Rb atoms, oriented on previously reported designs [50, 51, 52]. The resulting conductances are listed in table 2 alongside with respective DP dimensions.
Figure 9: Design of the differential pumping stages. Rubidium atoms are cooled in the 2D MOT at $x=0\,\mathrm{mm}$ and pushed toward the science cell with a divergence of 26 mrad (dotted red line, example from Ref. [51]). Lithium atoms leave the 2D MOT at $x=187\,\mathrm{mm}$ with a divergence of 50 mrad (dotted blue line, example from Ref. [29]) and are not further clipped by the Li differential pumping stage compared to the science cell opening angle (solid blue line). The maximum capturable divergence of rubidium and lithium atoms is therefore set by the opening angle of the Li DP and the science cell, respectively (solid lines).
@@ B.1 Numerical principle
The core of the simulations for the Zeeman slower beam is the numerical integration of the equations of motion obtained from the position and velocity-dependent radiation force acting on the atoms (cf. Eq. 3), using the python package scipy.intergrate version 1.15.0. The atoms’ starting position is the bottom of the oven, $190\text{\,}\mathrm{m}\mathrm{m}$ away from the 2D MOT center. The magnetic field is calculated with magpylib version 4.3.0, with the measured magnetization of $8.8\text{\times}{10}^{5}\text{\,}\mathrm{A}\mathrm{/}\mathrm{m}$ from [29].

### v2

@@ III.1 Evaluating the Li short-distance Zeeman slower
where $R$ is the loading rate, $\Gamma=1/\tau$ is the loss rate coefficient arising from background collisions, and $\beta$ is the two-body collisional loss coefficient. Solving this differential equation with the initial condition, $N_{0}=0$ yields the expression of the MOT dynamics (assuming isotropic density distribution)[dinneen1999]
@@
The Zeeman slower significantly boosts the atomic flux, thereby enhancing both the loading rate and the saturation atom number. Specifically, the loading rate increases by a factor of 44 to $R^{\mathrm{Li}}_{\mathrm{zs}}=6.639(25)\times 10^{9}~\mathrm{atoms/s}$ at a moderate oven temperature of $T_{\mathrm{Li}}=372~^{\circ}\mathrm{C}$. This temperature is $80~^{\circ}\mathrm{C}$ lower (23 times lower vapor pressure) than typical $^{6}\mathrm{Li}$ experiments using traditional Zeeman slowing techniques [wu2017LiK], which also contributes to prolonging the lifetime of the $^{6}\mathrm{Li}$ source.
@@ III.2 Li flux dependence
Figure 6: Optimization of the Zeeman slowing parameters. (a) 3D MOT loading rate as a function of Li oven temperature. The solid line represents the theoretical Li vapor pressure in the liquid phase [gehm2003]. (b) Loading rate dependence on the repumper-to-cooler intensity ratio. (c) Loading rate dependence on the Zeeman beam power. Measurements in panels (b) and (c) were conducted at a fixed oven temperature of $335\,^{\circ}\mathrm{C}$.
@@ IV Summary and Conclusion
During the preparation of this manuscript, we became aware of a compact high-flux ytterbium-rubidium Zeeman slower [chen2025].
@@ Appendix A Differential pumping channels design
It is essential to carefully design the dimensions of the differential pumping channels (DP) between the chambers to maintain a sufficient pressure ratio while not blocking the atomic flux more than necessary. The atomic beam divergence is given as its transverse-to-longitudinal velocity ratio, $\zeta=v_{\mathrm{tr}}/v_{x}$. The first is a result of the finite transverse temperature after the 2D MOT stage, while the latter can be tuned via the push beam power. In general, it is advantageous to reduce the distances from 2D MOTs to the science cell as much as possible, resulting in a larger opening angle of the science cell and thus an increased capturable fraction of atoms. The DP dimensions are then chosen to not clip the atomic beam significantly more than the science glass cell, which is shown schematically in Figure 9. The DPs consist of two sections in series (with diameters $d_{1}$ and $d_{2}$ as well as lengths $l_{1}$ and $l_{2}$), as it is easier to manufacture compared to one conical tube (as used, e.g., in [hammel2025, tobiashammel2021]). Additionally, the Rb DP is 45 polished on the surface facing the Rb atoms, oriented on previously reported designs [jollenbeck2011, holtkemeier2011, parry2015]. The resulting conductances are listed in table 2 alongside with respective DP dimensions.
Figure 9: Design of the differential pumping stages. Rubidium atoms are cooled in the 2D MOT at $x=0\,\mathrm{mm}$ and pushed toward the science cell with a divergence of 26 mrad (dotted red line, example from Ref. [holtkemeier2011]). Lithium atoms leave the 2D MOT at $x=187\,\mathrm{mm}$ with a divergence of 50 mrad (dotted blue line, example from Ref. [tiecke2009]) and are not further clipped by the Li differential pumping stage compared to the science cell opening angle (solid blue line). The maximum capturable divergence of rubidium and lithium atoms is therefore set by the opening angle of the Li DP and the science cell, respectively (solid lines).
@@ B.1 Numerical principle
The core of the simulations for the Zeeman slower beam is the numerical integration of the equations of motion obtained from the position and velocity-dependent radiation force acting on the atoms (cf. Eq. 3), using the python package scipy.intergrate version 1.15.0. The atoms’ starting position is the bottom of the oven, $190\text{\,}\mathrm{m}\mathrm{m}$ away from the 2D MOT center. The magnetic field is calculated with magpylib version 4.3.0, with the measured magnetization of $8.8\text{\times}{10}^{5}\text{\,}\mathrm{A}\mathrm{/}\mathrm{m}$ from [tiecke2009].